OpenStudy (haaileyyy):

Please Help Find The Product: (x+y)(x-7y)

OpenStudy (haaileyyy):

Would the answer be \[x ^{2}-7y ^{2}-7xy ^{2}\] ?

OpenStudy (mariamikayla):

(x+y)(x-7y)= x*x -x*7y+x*y-7y*y= =x^2 -7xy+xy-7y^2= =x^2-6xy-7y^2

OpenStudy (haaileyyy):

Thank You!

OpenStudy (mariamikayla):

You're welcome! :D